OpenStudy (anonymous):

Jacob has 10 rap, 18 rock, 8 country, and 4 pop CDs in his music collection. Two are selected at random. Find P(1 rap and 1 rock).

OpenStudy (anonymous):

I got you

OpenStudy (anonymous):

bro first you got to walk with me

OpenStudy (ybarrap):

You can get 1 rap and 1 rock in two ways E1 = rap on 1st rock on 2nd E2 = rock on 1st rap on 2nd $$ \text{P(1 rap and 1 rock)} = P(E1) + P(E2) \\ =\cfrac{10}{50}\times\cfrac{18}{49}+\cfrac{18}{50}\times\cfrac{10}{49}\\ =\cfrac{2\times 180}{50\times 49} $$